OpenStudy (anonymous):

the sides of a square are lengthened by 6cm, the area then becomes 196cm^2. what is the area of the original square? 172cm^2?

OpenStudy (anonymous):

Solve (x+6)^2=196

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

x^2 + 12x + 36 = 196

OpenStudy (anonymous):

Apply square root to both sides first is easier ag....go

OpenStudy (anonymous):

pixie! you been here all evening? dang. as mandolino said, solve \[(x+6)^2=198\] \[x+6=\sqrt{196}=14\] \[x=14-6=8\]

OpenStudy (anonymous):

so 32 is the area of the original square?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yeah, been here almost all night lol, trying to finish my final

OpenStudy (anonymous):

on line class?

OpenStudy (anonymous):

oh wait!!

OpenStudy (anonymous):

\[8^2=64\] not 32

OpenStudy (anonymous):

8*8=64

OpenStudy (anonymous):

so original square had area 64, not 32

OpenStudy (anonymous):

ok, so what would the length of one side be then? 8?

OpenStudy (anonymous):

online class yeah 17 more questions to go

OpenStudy (anonymous):

length of one side is 8, area is 64

OpenStudy (anonymous):

dang how many questions in this final, and how much time are you given?

OpenStudy (anonymous):

not due til tomorrow night, but still have another final to do, 2000+ word paper lol